OpenStudy (anonymous):

Find transfer function from below : a. ______s + 1___________ (s2+6s + 4) b. _______s2+3s+2________ (s+1)3 c. Find space node ( Tb ) from G(s).H(s) = ___________K_____________ + 0 (s+2) (s+4)

OpenStudy (anonymous):

? context?

OpenStudy (anonymous):

For a linear time-invariant system we can transform convolution into multiplication in the \(s\)-domain; whereas in the time domain we have $$y(t)=x(t)*h(t)$$we end up with $$Y(s)=X(s)H(s)$$ in the *complex frequency* domain; here, \(H(z)\) describes our frequency response and is thus known as our transfer function.

OpenStudy (anonymous):

What LTI system are we considering? is it just some standard sort-of filter?

OpenStudy (anonymous):

c. Find space node ( Tb ) from G(s).H(s) = ___________K_____________ = 0 (s+2) (s+4) i'm sorry , i wrong write sign

OpenStudy (anonymous):

what are (a), (b) asking for?

OpenStudy (anonymous):

i mind a and b just sign like a number..... i don't understand to transfer function from excercise what the relation with laplace methode

OpenStudy (anonymous):

plz sir, help me...

OpenStudy (anonymous):

I do not understand your question.

OpenStudy (anonymous):

Find Transfer function from : ________s+1__________ (s2+6s+4)

OpenStudy (anonymous):

That doesn't make sense though. That expression tells us nothing and could itself be a transfer function

OpenStudy (anonymous):

________s+1__________ (s^2+6s+4)

OpenStudy (anonymous):

Still does not make sense

OpenStudy (anonymous):

You are giving too little information

OpenStudy (anonymous):

how about this s^2 + 3s + 2 -------------- ( s + 1 ) ^3 still does not make sense, sir??